How many oxygen atoms are there in 95 g of barium chromate? Calculate the percent composition of oxygen by mass in How many oxygen atoms are there in 95g of barium chromate?
Answer #1

BaCrO4

Molar mass = 253 g / mol

Moles = mass / molar mass

= 95 / 253

= 0.3755

1 moles of BaCrO4 has 4 moles of oxygen atom look formula

So, moles of oxygen atom = 0.3755 * 4 = 1.502

No. Of oxygen atoms = mole * avagadro number

= 1.502 * 6.022*10^23

= 9.045*10^23

If you have any questions please comment

If you satisfied with the solution please rate it thanks